A portfolio consists of $14,400 in Stock M and $21,400 invested in Stock N. The expected return on these stocks is 8.50 percent and 12.10 percent, respectively. What is the expected return on the portfolio?
Answer #1

Total investment=(14400+21400)=35800

Portfolio return=Respective return*Respective investment weight

=(14400/35800*8.5)+(21400/35800*12.1)

which is equal to

=10.65%(Approx).